OpenStudy (anonymous):

Solve the equation 2 over 3 (x − 2) = 4x for x.

OpenStudy (anonymous):

impossibel because these are the only options I have negative 2 over 5 negative 5 over 2 2 over 5 5

OpenStudy (anonymous):

Hmm, maybe I'm reading the question wrong. Is it \[\frac{ 2 }{3x-2 } = 4x\]?

OpenStudy (anonymous):

2 ----- (x-2) = 4x 3

OpenStudy (anonymous):

Oh, I made a huge mistake. Sorry. 2(x-2) = 12x x-2 = 6x -2 = 5x x = -5/2

OpenStudy (anonymous):

then its 2(x-2)=4x*3 2x-4=12x 10x=-4 x=-4/10 x= negative 2 over 5 ....